OpenStudy (anonymous):

help Classify triangle BCD based on side lengths. (Points : 2) scalene isosceles equilateral right

OpenStudy (anonymous):

@jayz657

OpenStudy (anonymous):

its isoceles

OpenStudy (mayankdevnani):

@kendoll its isosceles because AC=BD AC=8 cm and BD=BE+ED=5+3=8 cm and those triangle whose two sides of same length is known as isosceles triangle.